# Helicity controls the direction of fluxes in rotating turbulence

### v1

@@ Abstract
Turbulence sustains out-of-equilibrium fluxes that are shaped by conservation laws. Three-dimensional flows conserve energy and sign-indefinite helicity, both being transferred to small scales. Here, we uncover a dual organization of energy fluxes in 3D rotating flows, shaped by helicity. When sufficiently-fast inertial waves interact with a large-scale 2D flow, they conserve their helicity separately by sign. This causes an inverse energy transfer, from 3D to 2D motions, which promotes self-organization and spectral condensation. In contrast, slower modes exchange helicity with modes of opposite helicity sign, similarly to non-rotating 3D turbulence. This generates a forward energy transfer, from the large-scale 2D flow to small 3D scales, coexisting with the inverse transfer. We determine analytically these bi-directional energy transfers to the 2D mean flow via a quasi-linear wave-kinetic theory. The theory captures the main Reynolds number and rotation rate dependence of the mean-flow amplitude in Navier-Stokes simulations from zero to infinite rotation rates.
@@
Inviscid invariants are expected to control out-of-equilibrium fluxes in turbulent systems. For example, in two dimensions, an inviscid fluid conserves energy and enstrophy (vorticity squared). These two sign-definite scale-related quantities are constrained to flow in opposite directions: energy to large scales and enstrophy to small scales [fjortoft1953changes, kraichnan1967inertial]. This leads to self-organization of turbulence into large-scale structures. In contrast, three-dimensional turbulence conserves energy and sign-indefinite helicity (the scalar product between velocity and vorticity) [kraichnan1973helical]. This leaves the spectral fluxes unconstrained, with a simultaneous transfer of both energy and helicity to small scales [chen2003joint, DitlevsenCascades, alexakis2017helically].
@@
Defying this classification, 3D fluids in the presence of external constraints, such as geometry, magnetic field or rotation, can exhibit a dual organization of energy, with simultaneous transfers to large and small scales [alexakis2018cascades, boffetta2011flux, smith1999transfer, falkovich2017vortices, rubio2014upscale, de2024pattern, benavides2017critical, van2019condensates, shaltiel2024direct, shavit2025turbulent]. Rotation, a primary ingredient in geophysical or astrophysical flows, with a crucial role in climate and weather modeling [vallis1993generation, ghil2020physics], provides a striking example: under rotation, three-dimensional fluctuations take the form of waves and self-organize into large-scale 2D structures [godeferd2015structure, shaltiel2024direct], although the invariants are identical to those in 3D turbulence.
@@
Meanwhile, when rotation is not too fast, 3D modes can also extract energy from large-scale 2D structures [seshasayanan2018condensates, clark2020phase], expected from, e.g., spontaneous generation of instabilities [sipp2000three, billant2021taylor]. Flows in the Earth’s oceans [balwada2022direct] and atmosphere [lindborg2005effect, czaja2019simulating] often lie in this regime. These bi-directional energy pathways in rotating 3D turbulence lack a general understanding: why are large scales energized in the absence of a sign-definite invariant? How can upscale and downscale energy fluxes coexist?
@@
Here we show that, in the presence of rotation, helicity crucially dictates the partition of energy. For moderate rotation rates, after a large-scale 2D flow spontaneously forms, two types of modes exist: fast waves, dominated by rotation, and slow waves, dominated by the 2D shear. For the former, 2D-3D interactions conserve helicity by sign so energy is transferred to the 2D flow, while the latter break this conservation law and extract energy from the 2D flow. We capture this behavior analytically via a mean-wave kinetic theory. Our work suggests a general mechanism for the generation of bi-directional fluxes in wave-dominated systems [alexakis2024large, labarre20242d].
@@
We consider the 3D Navier-Stokes equations (3DNSE) describing an incompressible fluid in a rotating frame, i.e. subject to a Coriolis force $2\Omega\mathbf{e}_{z}\times\bm{v}$, in a domain of dimensions $(L_{x},L_{y},L_{z})$. The flow is forced by a random three-dimensional isotropic forcing $\bm{f}$ where $\langle\bm{f}_{\bm{p}}(t)\bm{f}_{\bm{q}}(t^{\prime})\rangle=2\epsilon\chi_{\bm{p}}\delta(\bm{p}+\bm{q})\delta(t-t^{\prime})$, $\chi_{p}$ is restricted to a shell around $k_{f}=2\pi/l_{f}$ and energy is injected at rate $\epsilon$. The Reynolds number, $Re=\epsilon^{1/3}k_{f}^{-4/3}/\nu$, where $\nu$ is the kinematic viscosity, and the Rossby number, $Ro=\epsilon^{1/3}k_{f}^{2/3}/(2\Omega)$, jointly determine the resulting state. In the following we will also use the combined parameter $Ro_{\epsilon}=Ro\times\sqrt{Re}$. We use GHOST [mininni2011hybrid] to solve this system numerically, taking $l_{f}=L_{y}/10$ and $L_{x}=L_{z}=L_{y}/2=\pi$ (and including hyperviscosity).
@@
For low enough Ro, a large-scale 2D mean flow, termed a condensate, spontaneously emerges. This occurs after a transient where an inverse energy cascade among 2D modes develops once they are energized [alexakis2018cascades]. For our choice of domain geometry, the condensate takes the form of jets, see Fig. 1(a)-(b). We consider the statistical steady state with $\bm{U}=\langle\bm{v}\rangle=U(y)\mathbf{e}_{x}$ the jet mean flow. Typically, most of its energy is contained in the lowest mode $k_{y}=2\pi/L_{y}$, so we approximate $U(y)={U^{\prime}}\sqrt{2}/k_{y}~\sin(k_{y}y)$, where ${U^{\prime}}\equiv\sqrt{\int(\partial_{y}U)^{2}(dy/L_{y})}$ denotes the rms shear rate. Perhaps the most basic question is: what value does the condensate amplitude $U^{\prime}$ take as a function of $\text{Ro},\text{Re}$ and the box geometry? Here we answer this question in a mixed regime, where not all interactions with the condensate are wave-dominated, thereby completing [gome2025waves].
@@
The condensate is fueled by 3D waves [gome2025waves], called inertial waves. For each 3D mode $\bm{p}$ ($p_{z}\neq 0$) there are two wave polarizations $\bm{h}_{\bm{p}}^{s}$, such that $i\bm{p}\times\bm{h}_{\bm{p}}^{s}=sp\bm{h}_{\bm{p}}^{s}$, with chiralities $s=\pm 1$, dispersion relation $\omega_{\bm{p}}^{s}=s2\Omega p_{z}/p$, and with a given helicity sign $s=\pm 1$. $U^{\prime}$ is set by the energy transfer between the 3D fluctuations and the condensate, $T_{\rm 3D-2D}$, via the global energy balance:
@@
As a result of energy condensation, mean-wave interactions dominate over wave-wave interactions ($\bm{u}\cdot\nabla\bm{u}\ll\bm{U}\cdot\nabla\bm{u}$) for most scales (below a wavenumber denoted by $k_{U}$). Fluctuations are therefore governed by quasi-linear (QL) dynamics. Decomposing the fluctuations field into helical waves, $\bm{u}(\bm{x})=\sum_{\bm{p},s}a_{\bm{p}}^{s}\bm{h}_{\bm{p}}^{s}e^{i\omega_{\bm{p}}^{s}t+i\bm{p}\cdot\bm{x}}$, we obtain
$\hskip-10.00002pt\partial_{t}a_{\bm{p}}^{s}$ $=\hskip-10.00002pt\sum_{\begin{subarray}{c}\bm{q},\bm{k}=\pm\frac{2\pi}{L_{y}},\\ \tilde{s}=\pm s\end{subarray}}V_{\bm{p}\bm{k}\bm{q}}^{s\tilde{s}}U_{\bm{k}}^{*}a_{\bm{q}}^{\tilde{s}*}e^{-i(\omega_{\bm{p}}^{s}+\omega_{\bm{q}}^{\tilde{s}})t}\delta_{\bm{k}\bm{p}\bm{q}}+f_{\bm{p}}^{s}e^{-i\omega_{\bm{p}}^{s}t},$ (2)
with $\delta_{\bm{k}\bm{p}\bm{q}}=\delta_{\bm{k}+\bm{p}+\bm{q},0}$ and $V_{\bm{p}\bm{k}\bm{q}}^{s_{1},s_{2}}=-ip_{x}~(\bm{h}_{\bm{p}}^{-s_{1}}\cdot\bm{h}_{\bm{q}}^{-s_{2}})+ik_{y}~(h_{\bm{p},x}^{-s_{1}}~h_{\bm{q},y}^{-s_{2}})$ the coupling coefficients with the 2D mode.
A 3D mode $a_{\bm{p}}^{s}$ interacts with the condensate and with a companion mode $a_{\bm{q}}^{\tilde{s}}$ via Eq. 2. The interaction involves waves of either the same helicity sign $\tilde{s}=s$ (homochiral-wave interaction), or of opposite helicity signs $\tilde{s}=-s$ (heterochiral-wave interaction). In addition, the helicity exchange between the large-scale 2D flow and the fluctuations is negligible due to scale separation $l_{f}/L_{y}\ll 1$, so the helicity of the fluctuations is conserved separately — see [gome2025waves] and Supplementary Material (SM). Species with opposite helicites can therefore exchange energy (and helicity) only through heterochiral-wave interactions with the condensate.
In classical 3D turbulence, the exchange of energy and helicity between opposite helicity modes [kraichnan1973helical] results in the transfer of both invariants to smaller scales [chen2003joint, alexakis2017helically]. As shown in Fig. 1(d), and explained throughout this Letter, introducing rotation progressively suppresses heterochiral-wave interactions, with a profound impact on the energy fluxes. Measuring the energy contribution from waves of either the same helicity sign (hollow circles), or of opposite helicity signs (colored circles), we observe that the condensate is fueled by the former, while the latter extract energy from it.
@@
$|\omega_{\bm{p}}^{s}+\omega_{\bm{q}}^{\tilde{s}}|<{U^{\prime}},~~~\omega_{\bm{p}}^{s}=s2\Omega\frac{p_{z}}{p},$ (3)
for which the frequency sum $\omega_{\bm{p}}^{s}+\omega_{\bm{q}}^{\tilde{s}}$ is sufficiently small so that the interaction does not average-out over the slow time scale $t=o(1/{U^{\prime}})$. This can be shown by following wave kinetic theory [zakharov2012kolmogorov, nazarenko2011wave, gome2025waves], considering the dynamics at the lowest order in time-scale separation ${U^{\prime}}/\Omega$ [gome2025waves]. The larger the rotation, the more restrictive condition (3) becomes, decreasing the number of possible interactions.
Condition (3) is more easily satisfied by homochiral-wave interactions ($\tilde{s}=s$), for which it reads $|\omega_{p}^{s}+\omega_{q}^{s}|\approx|k_{y}\partial_{p}\omega_{p}^{s}|<U^{\prime}$ assuming scale separation $l_{f}/L_{y}\ll 1$, than for heterochiral-wave interactions, where it requires $\omega_{\bm{p}}^{+}+\omega_{\bm{q}}^{-}\approx 2\omega_{\bm{p}}^{+}<U^{\prime}$. Here we work in the regime where homochiral-wave interactions fulfill resonant condition (3), $Ro_{\epsilon}>l_{f}/(2L_{y})$ (see [gome2025waves]).
For $\tilde{s}=-s$, a mode $\bm{p}$ is involved only in homochiral-wave interactions if it violates the resonance condition (3), and participates in both types of interactions if it satisfies it. Mode $\bm{p}$ therefore belongs to one of two sectors:
@@
illustrated in Fig. 2(a). Following (4), modes in sector H have fast frequencies, corresponding to a wave-dominated behavior. In contrast, the dynamics in sector A is shear-dominated and essentially unaffected by rotation. As the frequency $\omega_{\bm{p}}^{s}$ decreases with $p$, a transition from sector H to sector A occurs when $p>p^{*}\equiv\frac{4\Omega|p_{z}|}{{U^{\prime}}}$. The rotation rate, via $U^{\prime}/\Omega$, therefore controls the number of modes in sectors $A$ and $H$.
The mean flow $U(y)\mathbf{e}_{x}$ tends to shear the waves, moving their energy to increasing values of $|p_{y}|$, while leaving $p_{x}$ and $p_{z}$ unchanged. This generates a positive energy flux towards large $|p_{y}|$, denoted by $\Pi^{s}_{\bm{p}}$ for waves of helicity sign $s$, whose magnitude at small scales remains to be determined. Energy injected in Sector H thus flows into Sector A ($p>p^{*})$, see Fig. 2(b). In the following we will compute the energy flux $\Pi^{s}_{\bm{p}}$ in sectors H and A separately. This flux will, in turn, determine the energy transfer to the condensate from modes with given ($p_{x},p_{z},s$), equal to the difference between the energy injected into these modes, $\epsilon^{s}_{p_{x},p_{z}}$, and the flux to arbitrarily small scales: $\epsilon^{s}_{p_{x},p_{z}}-\Pi^{s}_{|p_{y}|\to\infty}$.
The spectral energy flux $\Pi_{\bm{p}}^{s}$ is constrained by the conservation of single-sign helicity $H_{\bm{p}}^{s}=sp\langle|a_{\bm{p}}^{s}|^{2}\rangle/2$, occurring in sector H but not in A. This can be seen from the spectral balance for the wave helicity, obtained using scale separation in the QL system:
$\partial_{t}H_{\bm{p}}^{s}+\partial_{p_{y}}(sp\Pi^{s}_{\bm{p}})$ $=-sp~\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\bm{p}\bm{k}}^{s,-s}+sp\epsilon\chi_{\bm{p}}^{s},$ (5)
where $\Pi^{s}_{\bm{p}}=-{U^{\prime}}(\langle a_{\bm{p}}^{s}a_{-\bm{p}+\bm{k}}^{s}\rangle+\langle a_{\bm{p}}^{s}a_{-\bm{p}-\bm{k}}^{s}\rangle)/(2\sqrt{2})$ and $\langle uv\rangle_{\bm{p}\bm{k}}^{s,-s}=\Re\left[h_{\bm{p},x}^{s}h_{-\bm{p},y}^{-s}\left(\langle a_{\bm{p}}^{s}a_{-\bm{p}+\bm{k}}^{-s}\rangle+\langle a_{\bm{p}}^{s}a_{-\bm{p}-\bm{k}}^{-s}\rangle\right)\right]$ is the Reynolds stress related to waves of opposite chiralities. The transfer of helicity (and energy) between species $s$ and s-s is driven by the opposite-chirality correlators $\langle a_{\bm{p}}^{s}a_{-\bm{p}+\bm{k}}^{-s}\rangle$ entering $\langle uv\rangle_{\bm{p}\bm{k}}^{s,-s}$ in Eq. (5).

### v2

@@ Abstract
Turbulence sustains out-of-equilibrium energy fluxes shaped by conservation laws. Three-dimensional flows conserve energy and sign-indefinite helicity, both being transferred to small scales. Yet in 3D rotating turbulence, energy is observed to flow simultaneously toward large-scale two-dimensional structures and toward small-scale three-dimensional waves. We uncover the origin of this dual behavior. When sufficiently fast inertial waves interact with a large-scale 2D flow, they conserve their helicity separately by sign, enforcing an inverse transfer of energy from 3D waves to 2D motions and promoting spectral condensation. Slower modes, by contrast, exchange helicity across opposite-sign sectors and thus behave as in non-rotating turbulence, driving a forward transfer from the large-scale 2D flow to small 3D scales. Using a mean–wave kinetic theory, we derive analytical expressions for these competing bi-directional transfers and quantitatively predict the rotation- and Reynolds-number dependence of the large-scale 2D flow in fully nonlinear simulations, unifying the picture from zero to infinite rotation.
@@
Inviscid invariants are expected to control out-of-equilibrium fluxes in turbulent systems. For example, in two dimensions, an inviscid fluid conserves energy and enstrophy (vorticity squared). These two sign-definite scale-related quantities are constrained to flow in opposite directions: energy to large scales and enstrophy to small scales [21, 29]. This leads to self-organization of turbulence into large-scale structures. In contrast, three-dimensional turbulence conserves energy and sign-indefinite helicity (the scalar product between velocity and vorticity) [30]. This leaves the spectral fluxes unconstrained, with a simultaneous transfer of both energy and helicity to small scales [11, 17, 4].
@@
Defying this classification, 3D fluids in the presence of external constraints, such as geometry, magnetic field or rotation, can exhibit a dual organization of energy, with simultaneous transfers to large and small scales [1, 9, 49, 19, 41, 15, 6, 51, 45, 47]. Rotation, a primary ingredient in geophysical or astrophysical flows, with a crucial role in climate and weather modeling [50, 24], provides a striking example: under rotation, three-dimensional fluctuations take the form of waves and self-organize into large-scale 2D structures [25, 45], although the invariants are identical to those in 3D turbulence.
@@
Meanwhile, when rotation is not too fast, 3D modes can also extract energy from large-scale 2D structures [43, 13], expected from, e.g., spontaneous generation of instabilities [48, 8]. Flows in the Earth’s oceans [5] and atmosphere [35, 14] often lie in this regime. These bi-directional energy pathways in rotating 3D turbulence lack a general understanding: why are large scales energized in the absence of a sign-definite invariant? How can upscale and downscale energy fluxes coexist?
@@
Here we show that, in the presence of rotation, helicity crucially dictates the partition of energy. For moderate rotation rates, after a large-scale 2D flow spontaneously forms, two types of modes exist: fast waves, dominated by rotation, and slow waves, dominated by the 2D shear. For the former, 2D-3D interactions conserve helicity by sign so energy is transferred to the 2D flow, while the latter break this conservation law and extract energy from the 2D flow. We capture this behavior analytically via a mean-wave kinetic theory, validated with numerical simulations. Our work suggests a general mechanism for the generation of bi-directional fluxes in wave-dominated systems [2, 32].
@@
We consider the 3D Navier-Stokes equations (3DNSE) describing an incompressible fluid in a rotating frame, i.e. subject to a Coriolis force $2\Omega\mathbf{e}_{z}\times\boldsymbol{v}$, in a domain of dimensions $(L_{x},L_{y},L_{z})$. The flow is forced by a random three-dimensional isotropic forcing $\boldsymbol{f}$ where $\langle f_{\boldsymbol{p},i}(t)f_{\boldsymbol{q},j}(t^{\prime})\rangle=2\epsilon(1-p_{i}p_{j}/p^{2})\chi_{\boldsymbol{p}}\delta(\boldsymbol{p}+\boldsymbol{q})\delta(t-t^{\prime})$, $\chi_{p}$ is restricted to a shell around $k_{f}=2\pi/l_{f}$ and energy is injected at rate $\epsilon$. The Reynolds number, $Re=\epsilon^{1/3}k_{f}^{-4/3}/\nu$, where $\nu$ is the kinematic viscosity, and the Rossby number, $Ro=\epsilon^{1/3}k_{f}^{2/3}/(2\Omega)$, jointly determine the resulting state. In the following we will also use the combined parameter $Ro_{\epsilon}=Ro\times\sqrt{Re}$. We use GHOST [38] to solve this system numerically, taking $l_{f}=L_{y}/10$ and $L_{x}=L_{z}=L_{y}/2=\pi$ (and including hyperviscosity).
@@
For low enough Ro, a large-scale 2D mean flow, termed a condensate, spontaneously emerges. This occurs after a transient where an inverse energy cascade among 2D modes develops after they are energized [1]. For our choice of domain geometry, the condensate takes the form of jets, see Fig. 1(a)-(b). We consider the statistical steady state with $\boldsymbol{U}=\langle\bm{v}\rangle=U(y)\mathbf{e}_{x}$ the jet mean flow. Typically, most of its energy is contained in the lowest mode $k_{y}=2\pi/L_{y}$, so we approximate $U(y)={U^{\prime}}\sqrt{2}/k_{y}~\sin(k_{y}y)$, where ${U^{\prime}}\equiv\sqrt{\int(\partial_{y}U)^{2}(dy/L_{y})}$ denotes the rms shear rate. Perhaps the most basic question is: what value does the condensate amplitude $U^{\prime}$ take as a function of $\text{Ro},\text{Re}$ and the box geometry? Here we answer this question in a mixed regime, where not all interactions with the condensate are wave-dominated, thereby completing [26].
@@
The condensate is fueled by 3D waves [26], called inertial waves. For each 3D mode $\boldsymbol{p}$ ($p_{z}\neq 0$) there are two wave polarizations $\boldsymbol{h}_{\boldsymbol{p}}^{s}$, such that $i\boldsymbol{p}\times\boldsymbol{h}_{\boldsymbol{p}}^{s}=sp\boldsymbol{h}_{\boldsymbol{p}}^{s}$, with chiralities $s=\pm 1$, dispersion relation $\omega_{\boldsymbol{p}}^{s}=s2\Omega p_{z}/p$, and with a given helicity sign $s=\pm 1$. $U^{\prime}$ is set by the energy transfer between these 3D fluctuations and the condensate, $T_{\rm 3D-2D}$, via the global energy balance:
@@
As a result of energy condensation, mean-wave interactions dominate over wave-wave interactions ($\boldsymbol{u}\cdot\nabla\boldsymbol{u}\ll\boldsymbol{U}\cdot\nabla\boldsymbol{u}$) for most scales (below a wavenumber denoted by $k_{U}$). Fluctuations are therefore governed by quasi-linear (QL) dynamics. Decomposing the fluctuations field into helical waves, $\boldsymbol{u}(\boldsymbol{x})=\sum_{\boldsymbol{p},s}a_{\boldsymbol{p}}^{s}\boldsymbol{h}_{\boldsymbol{p}}^{s}e^{i\omega_{\boldsymbol{p}}^{s}t+i\boldsymbol{p}\cdot\boldsymbol{x}}$, we obtain
$\hskip-10.00002pt\partial_{t}a_{\boldsymbol{p}}^{s}$ $=\hskip-10.00002pt\sum_{\begin{subarray}{c}\boldsymbol{q},\boldsymbol{k}=\pm\frac{2\pi}{L_{y}},\\ \tilde{s}=\pm s\end{subarray}}V_{\boldsymbol{p}\boldsymbol{k}\boldsymbol{q}}^{s\tilde{s}}U_{\boldsymbol{k}}^{*}a_{\boldsymbol{q}}^{\tilde{s}*}e^{-i(\omega_{\boldsymbol{p}}^{s}+\omega_{\boldsymbol{q}}^{\tilde{s}})t}\delta_{\boldsymbol{k}\boldsymbol{p}\boldsymbol{q}}+f_{\boldsymbol{p}}^{s}e^{-i\omega_{\boldsymbol{p}}^{s}t},$ (2)
with $\delta_{\boldsymbol{k}\boldsymbol{p}\boldsymbol{q}}=\delta_{\boldsymbol{k}+\boldsymbol{p}+\boldsymbol{q},0}$ and $V_{\boldsymbol{p}\boldsymbol{k}\boldsymbol{q}}^{s_{1},s_{2}}=-ip_{x}~(\boldsymbol{h}_{\boldsymbol{p}}^{-s_{1}}\cdot\boldsymbol{h}_{\boldsymbol{q}}^{-s_{2}})+ik_{y}~(h_{\boldsymbol{p},x}^{-s_{1}}~h_{\boldsymbol{q},y}^{-s_{2}})$ the coupling coefficients with the 2D mode.
A 3D mode $a_{\boldsymbol{p}}^{s}$ interacts with the condensate and with a companion mode $a_{\boldsymbol{q}}^{\tilde{s}}$ via Eq. 2. The interaction involves waves of either the same helicity sign $\tilde{s}=s$ (homochiral-wave interaction), or of opposite helicity signs $\tilde{s}=-s$ (heterochiral-wave interaction). In addition, the helicity exchange between the large-scale 2D flow and the fluctuations is negligible due to scale separation $l_{f}/L_{y}\ll 1$, so the helicity of the fluctuations is conserved separately — see [26] and Supplementary Material (SM). Species with opposite helicites can therefore exchange energy (and helicity) only through heterochiral-wave interactions with the condensate.
In classical 3D turbulence, the exchange of energy and helicity between opposite helicity modes [30] results in the transfer of both invariants to smaller scales [11, 4]. As shown in Fig. 1(d), and explained throughout this Letter, introducing rotation progressively suppresses heterochiral-wave interactions, with a profound impact on the energy fluxes. Measuring the energy contribution from waves of either the same helicity sign (hollow circles), or of opposite helicity signs (colored circles), we observe that the condensate is fueled by the former, while the latter extract energy from it.
@@
$|\omega_{\boldsymbol{p}}^{s}+\omega_{-\boldsymbol{p}-\boldsymbol{k}}^{\tilde{s}}|<{U^{\prime}},~~~\omega_{\boldsymbol{p}}^{s}=s2\Omega\frac{p_{z}}{p},$ (3)
for which the frequency sum $\omega_{\boldsymbol{p}}^{s}+\omega_{-\boldsymbol{p}-\boldsymbol{k}}^{\tilde{s}}$ is sufficiently small so that the interaction does not average-out over the slow time scale $t=o(1/{U^{\prime}})$. This can be shown by following wave kinetic theory [52, 39, 26], considering the dynamics at the lowest order in time-scale separation ${U^{\prime}}/\Omega$ [26]. The larger the rotation, the more restrictive condition (3) becomes, decreasing the number of possible interactions.
Condition (3) is more easily satisfied by homochiral-wave interactions ($\tilde{s}=s$), for which it reads $|\omega_{\boldsymbol{p}}^{s}+\omega_{-\boldsymbol{p}-\boldsymbol{k}}^{s}|\approx|k_{y}\partial_{p}\omega_{\boldsymbol{p}}^{s}|<U^{\prime}$ assuming scale separation $l_{f}/L_{y}\ll 1$, than for heterochiral-wave interactions, where it requires $\omega_{\boldsymbol{p}}^{+}+\omega_{-\boldsymbol{p}-\boldsymbol{k}}^{-}\approx 2\omega_{\boldsymbol{p}}^{+}<U^{\prime}$. Here we work in the regime where homochiral-wave interactions fulfill resonant condition (3), $Ro_{\epsilon}>l_{f}/(2L_{y})$ (see [26]).
For $\tilde{s}=-s$, a mode $\boldsymbol{p}$ is involved only in homochiral-wave interactions if it violates the resonance condition (3), and participates in both types of interactions if it satisfies it. Mode $\boldsymbol{p}$ therefore belongs to one of two sectors:
@@
illustrated in Fig. 2(a). Following (4), modes in sector H have fast frequencies, corresponding to a wave-dominated behavior. In contrast, the dynamics in sector A is shear-dominated and essentially unaffected by rotation. As the frequency $\omega_{\boldsymbol{p}}^{s}$ decreases with $p$, a transition from sector H to sector A occurs when $p>p^{*}\equiv\frac{4\Omega|p_{z}|}{{U^{\prime}}}$. The rotation rate, via $U^{\prime}/\Omega$, therefore controls the number of modes in sectors $A$ and $H$.
The mean flow $U(y)\mathbf{e}_{x}$ tends to shear the waves, moving their energy to increasing values of $|p_{y}|$, while leaving $p_{x}$ and $p_{z}$ unchanged. This generates a positive energy flux towards large $|p_{y}|$, denoted by $\Pi^{s}_{\bm{p}}$ for waves of helicity sign $s$, whose magnitude at small scales remains to be determined. Energy injected in Sector H thus flows into Sector A ($p>p^{*})$, see Fig. 2(b). In the following we will compute the energy flux $\Pi^{s}_{\boldsymbol{p}}$ in sectors H and A separately. This flux will, in turn, determine the energy transfer to the condensate from modes with given ($p_{x},p_{z},s$), equal to the difference between the energy injected into these modes, $\epsilon^{s}_{p_{x},p_{z}}$, and the flux to arbitrarily small scales: $\epsilon^{s}_{p_{x},p_{z}}-\Pi^{s}_{|p_{y}|\to\infty}$.
The spectral energy flux $\Pi_{\boldsymbol{p}}^{s}$ is constrained by the conservation of single-sign helicity $H_{\boldsymbol{p}}^{s}=sp\langle|a_{\boldsymbol{p}}^{s}|^{2}\rangle/2$, occurring in sector H but not in A. This can be seen from the spectral balance for the wave helicity, obtained using scale separation in the QL system:
$\partial_{t}H_{\boldsymbol{p}}^{s}+\partial_{p_{y}}(sp\Pi^{s}_{\boldsymbol{p}})$ $=-sp~\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{s,-s}+sp\epsilon\chi_{\boldsymbol{p}}^{s},$ (5)
where $\Pi^{s}_{\boldsymbol{p}}=-{U^{\prime}}(\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}+\boldsymbol{k}}^{s}\rangle+\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}-\boldsymbol{k}}^{s}\rangle)/(2\sqrt{2})$ and $\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{s,-s}=\Re\left[h_{\boldsymbol{p},x}^{s}h_{-\boldsymbol{p},y}^{-s}\left(\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}+\boldsymbol{k}}^{-s}\rangle+\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}-\boldsymbol{k}}^{-s}\rangle\right)\right]$ is the Reynolds stress related to waves of opposite chiralities. The transfer of helicity (and energy) between species $s$ and s-s is driven by the opposite-chirality correlators $\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}+\boldsymbol{k}}^{-s}\rangle$ entering $\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{s,-s}$ in Eq. (5).